OpenStudy (anonymous):

a=1/2bh how do you solve for b?

OpenStudy (anonymous):

a = 1/2bh Solve for b. To solve for a variable means to isolate the variable on one side of the equation. NOTE: We can write (1/2)(bh) as one fraction like this (bh/2). They mean the same thing. We now have this: a = (bh/2) We multiply both sides of the fractional equation by 2. This is done to remove the fraction. Doing this we now have a linear equation: 2a = bh We now divide both sides of the equation by h to find the value of b. Final answer: (2a/h) = b Did you follow?